OpenStudy (anonymous):

Find the product. (fg)^9

OpenStudy (turingtest):

are you given definitions for f ang g? or is this ALL the info that you are given in the question?

OpenStudy (anonymous):

all the info

OpenStudy (turingtest):

then just use the distributive law of exponents\[\large(xy)^a=x^a\cdot y^a\]